Suppose that $27.40 is withheld for federal income tax and $12.35 is
withheld for Social Security. What would your net pay be? _____
SALES CLERK
Monday - Friday
8-hr. day $4.35/hr.
Apply in person.
Pedal Power Bike Shop

Answer by rfer(16322)

8*5=40 hr per wk
40*4.35=174.00
174.00-27.40-12.35=$134.25 net